find the slope and y-intercept of the line in the given equation y=-4/5x-10
Answer by mananth(16946)

find the slope and y-intercept of the line in the given equation y=-4/5x-10
...
y=-4/5x-10
compare with equation
y = mx +b
slope =m=(-4/5)
y intercept =b= -10
...
m.ananth@hotmail.ca